# A Dataset and Benchmark for Hospital Course Summarization with Adapted Large Language Models

### v2

@@ CONCLUSION
Our work spans multiple elements of data assimilation [albers2020assimilation]. Our methods utilize clinical (1) data sources like MIMIC-IV [johnson2023mimic] to integrate (2) data translation, where we preprocess raw clinical notes to standardize them for machine learning tasks, and (3) knowledge creation, where we adapt LLMs for BHC synthesis. By demonstrating the effectiveness of LLMs through a comprehensive evaluation strategy, we pave the way for their future (4) deployment in clinical decision-making scenarios.
@@
AUTHOR CONTRIBUTIONS
AA designed and executed the entire study, led the acquisition and development of the novel dataset, and drafted the manuscript, also undertaking the knowledge creation element from data assimilation [albers2020assimilation]. DVV provided essential methodological expertise, contributed to experimental design, and assisted in data analysis and manuscript writing. YIA contributed to the clinical study analysis and assisted in manuscript preparation. JH, CB, EPR, SG, and NC provided clinical expertise through a review of clinical notes with their respective summaries and contributed to the manuscript review. JD, AST, and JK contributed to dataset development, LLM pipeline establishment, and provided critical feedback on the manuscript. ASC oversaw the project and contributed to study design, data interpretation, and manuscript writing. All authors contributed to a critical review of the manuscript and approved the final version.
ACKNOWLEDGEMENTS
Computing resources were partially provided by One Medical and Stanford University. Microsoft provided Azure OpenAI credits for this project via the Accelerate Foundation Models Academic Research (AFMAR) program. A.C. receives research support from NIH grants R01 HL167974, R01 HL169345, R01 AR077604, R01 EB002524, R01 AR079431, and P41 EB027060; from NIH contracts 75N92020C00008 and 75N92020C00021; from Stanford Center for Artificial Intelligence and Medicine, Stanford Institute for Human Centered AI, from Stanford Center for Digital Health, from Stanford Cardiovascular Institute, from Stanford Center for Precision Health and Integrated Diagnostics; from GE Healthcare, Philips and Amazon.
COMPETING INTERESTS
@@ ACKNOWLEDGEMENTS
DATA AVAILABILITY
The MIMIC-IV-BHC [aali2024mimic] dataset is currently under review for publication on PhysioNet. We curated this dataset from the original MIMIC-IV-Note [johnson2023mimic] dataset, also published on PhysioNet.
CODE AVAILABILITY

### v3

@@ CONCLUSION
CONTRIBUTORSHIP STATEMENT
AA designed and executed the entire study, led the acquisition and development of the novel dataset, and drafted the manuscript, effectively undertaking the knowledge creation element from data assimilation [albers2020assimilation]. DVV provided essential methodological expertise, contributed to experimental design, and assisted in data analysis and manuscript writing. YIA contributed to the clinical study analysis and assisted in manuscript preparation. JH, CB, EPR, SG, and NC provided clinical expertise through a review of clinical notes with their respective summaries and contributed to the manuscript review. JD, AST, and JK contributed to dataset development, LLM pipeline establishment, and provided critical feedback on the manuscript. ASC oversaw the project and contributed to study design, data interpretation, and manuscript writing. All authors contributed to a critical review of the manuscript and approved the final version.
FUNDING STATEMENT
A.C. receives research support from NIH grants R01 HL167974, R01 HL169345, R01 AR077604, R01 EB002524, R01 AR079431, and P41 EB027060; from NIH contracts 75N92020C00008 and 75N92020C00021; from Stanford Center for Artificial Intelligence and Medicine, Stanford Institute for Human Centered AI, from Stanford Center for Digital Health, from Stanford Cardiovascular Institute, from Stanford Center for Precision Health and Integrated Diagnostics; from GE Healthcare, Philips and Amazon. Computing resources were partially provided by One Medical and Stanford University. Microsoft provided Azure OpenAI credits for this project via the Accelerate Foundation Models Academic Research (AFMAR) program.
DATA AVAILABILITY STATEMENT
To allow researchers to replicate our adaptation and evaluation approach for BHC summarization, we publicly release:
  1. 1.
    The MIMIC-IV-BHC dataset, published on PhysioNet [aali2024mimic].
  2. 2.
    The underlying code for data pre-processing, LLM adaptation, and evaluation: github.com/StanfordMIMI/clin-bhc-summ.
For this study, we use Microsoft Azure API versions GPT-3.5-Turbo-0613 and GPT-4-0613, as permitted under the guidelines set by PhysioNet, following all dataset usage protocols. Furthermore, this study was institutional review board exempt because it did not meet the criteria for human subjects research. Due to the use of MIMIC data in fine-tuning, we are unable to publicly share the fine-tuned model weights. However, these weights are available to credentialed users of MIMIC datasets upon request.
COMPETING INTERESTS STATEMENT